OpenStudy (taylor.joe):

Does anyone know how to do this???? Using the properties of exponents, apply the rational exponent to the numerator and the denominator, and then rationalize the denominator. C= square root of (E/M)

OpenStudy (fibonaccichick666):

are you given any values?

OpenStudy (taylor.joe):

Nope. That's all the give me. Which is why i don't understand.

rebeccaxhawaii (rebeccaxhawaii):

^^

OpenStudy (fibonaccichick666):

ok, so, let's try a few cases first so you can generalize. It'll make more sense

OpenStudy (fibonaccichick666):

so let's say you had \[\sqrt{\frac{4}{9}}\] How would you find that?

OpenStudy (taylor.joe):

Ummm I'm not exactly sure.

OpenStudy (taylor.joe):

Sorry im really slow when it comes to understanding in math.

OpenStudy (fibonaccichick666):

Did you miss days? Cause before you get into your problem, you should know how to do this

OpenStudy (taylor.joe):

Actually, hate to admit this but i skipped around and karmas kicking me in the butt for it. I've looked it up but i still don't understand.

OpenStudy (fibonaccichick666):

ok, so. watch this video first, then I can help you https://www.khanacademy.org/math/algebra-basics/core-algebra-foundations/square-roots-for-college/v/square-roots-and-real-numbers

OpenStudy (fibonaccichick666):

At like 6:55 you will find a fraction problem he explains

OpenStudy (taylor.joe):

Okay. Thank you! I really appreciate the help.

OpenStudy (fibonaccichick666):

np, so after you watch that, try my square root problem above. I will be back on when I get home. That's about 30 minutes

OpenStudy (taylor.joe):

Awesome! Thanks.

OpenStudy (taylor.joe):

\[\sqrt{\frac{ 4 }{ 9 }}\] That would be \[\frac{ 2 }{ 3 }\] right?

OpenStudy (taylor.joe):

I got it now! Thanks for your help. Your helped allot!!! :-)

OpenStudy (fibonaccichick666):

yea, that's right. So now, do you get how to do yours?

OpenStudy (taylor.joe):

Yes. I'm pretty sure. If not then i'm kinda screwed because i already turned it in. Thank you so much. It really helped.

OpenStudy (fibonaccichick666):

oh ok, well, best of luck to you then!

OpenStudy (taylor.joe):

Thanks!!!!!